OpenStudy (anonymous):

If you draw three cards from a regular deck of 52 cards, keeping each card out of the deck after you draw it, what are the chances you draw a seven and two queens?

OpenStudy (anonymous):

The probability you draw a 7 first and then 2 queens is: 4/52 x 4/51 x 3/50 = 2/5525 You need to multiply that by 3 though since you might get Queen... then a 7...then another Queen or Queen ... Queen... then 7 3 x 2/5525 = 6/5525 or 0.001086